The highest energy a particle emitted in the decay of 238U to 234Th is 

4196 &- 4 keV. From this information and the known mass of 238U, compute 

the mass of 234Th


1
Expert's answer
2021-06-13T17:35:42-0400
"m_{Th}=m_{U}-m_{He}-\frac{E}{c^2}\\m_{Th}=238.05-4.00-\frac{4.274}{931.5}\\
m_{Th}=234.04\ u"